A rectangle is 10 mm. longer than its width and its perimeter is 52 mm. Find the dimensions of the rectangle.
Thanks a lot if anyone could help me solve this problem.
Beth

Answer by askmemath(368)

Hello Beth
Let the width = W
Then length = W+10
Perimeter of a rectanlge = 2x(l+w)= 52
52= 2(w+w+10)
Dividing by 2 throughout
26 = 2w+10
Subtracting 10 from both sides
16 = 2w
Again dividing by 2 throughout
8 = w
which means L = 8+10 = 18